# Yong Liu    Guo Qin    Zhiyuan Shi    Zhi Chen    Caiyin Yang    Xiangdong Huang    Jianmin Wang    Mingsheng Long

### v2

@@ C.2 Scaling Behavior Using More Data
We compare Sundial with other time series foundation models that are pre-trained with smaller datasets: Chronos (Ansari et al., 2024) is pre-trained on 94 billion time points, and Moirai is pre-trained on 230 billion time points. As their pre-training datasets are part of the subset of TimeBench, we also conduct pre-training on Sundial using these subsets. As shown in Table 8, these results highlight the scaling behavior of Sundial using larger datasets. Additionally, Sundial still achieves better zero-shot forecasting than its counterpart models with the same pre-training dataset.
Table 8: Zero-shot forecasting performance of models trained on different scales of datasets (measured in time points, pts, and 1B means a billion). We report the averaged results from four prediction lengths $\{96,192,336,720\}$ on Time-Series-Library (Wu et al., 2022).
@@ C.3 Performance with Varying Lookback Lengths
Figure 10: Zero-shot forecasting performance using different lookback lengths in $\{480,960,1440,1920,2400,2880\}$. We report the averaged results from four prediction lengths $\{96,192,336,720\}$ on Time-Series-Library (Wu et al., 2022).
@@ C.4 Zero-Shot Results of Point Forecasting
Table 9 provides full zero-shot results on Time-Series-Library forecasting benchmark (Wu et al., 2022), including prediction horizons in $\{96,192,336,720\}$. We build Sundial with different model sizes with configurations in Table 5. The context length is fixed as $2880$. We truncate the model’s predictions for tasks requiring a prediction length less than $F=720$.
We compare the most advanced time series foundation models based on their official checkpoints, including Time-MoE (Shi et al., 2024b), Timer (Liu et al., 2024a, b), Moirai (Woo et al., 2024), TimesFM (Das et al., 2023b), and Chronos (Ansari et al., 2024). We conduct zero-shot evaluations on datasets that are not included during the pre-training of the corresponding models. For each of the evaluated model, we use their maximum input length during inference. Metrics (MSE/MAE) are calculated from all predicted windows in the test split of each dataset following Liu et al. (2024a).
@@ C.5 Zero-Shot Results on GIFT-Eval and FEV Leaderboard
We evaluate our models on GIFT-Eval, a benchmark designed to comprehensively assess forecasting performance across diverse time series. GIFT-Eval includes $23$ datasets covering $144,000$ time series and $177$ million data points, which constitute a total of $97$ forecasting configurations. We use the official evaluation suite established by the research team of Salesforce and report aggregated results in Table 2. We evaluate the performance and inference time on the FEV leaderboard, which was originally proposed by Ansari et al. (2024) and established by AutoGluon, which comprises $27$ datasets for zero-shot evaluation. We report aggregated metrics in Figure 4 and assess the inference time in Figure 5. We released the detailed results by submitting Sundial to their open benchmark [^2] https://huggingface.co/spaces/Salesforce/GIFT-Eval..
@@ D.1 Showcases of Sundial
Figure 11-13 present zero-shot forecasting showcases on all the datasets from FEV (Ansari et al., 2024) and TSLib (Wu et al., 2022). By generating $20$ predictions with different initial noise, we estimate the median and $80\%$ prediction interval.
Table 9: Zero-shot forecasting results of time series foundation models on long-term forecasting datasets (Wu et al., 2022). A lower MSE or MAE indicates a better prediction. Averaged results of four prediction lengths are reported here. $1^{\text{st}}$ Count represents the number of wins achieved by a model under all prediction lengths and datasets. Results of baseline models are officially reported by Shi et al. (2024b). Datasets for pre-training are not evaluated on corresponding models, which are denoted by the dash (--).
@@
[…]
@@
  • $\ast$ Traffic (PEMS, ) is not evaluated because it is included in the pre-training datasets of these time series foundation models.
Figure 11: Showcases of zero-shot predictions from Sundial (Base) on the FEV leaderboard (Ansari et al., 2024).
Figure 12: Showcases of zero-shot predictions from Sundial (Base) on the FEV leaderboard (Ansari et al., 2024).
Figure 13: Showcases of zero-shot predictions from Sundial (Base) on long-term forecasting datasets (Wu et al., 2022).

### v3

@@ C.2 Scaling Behavior Using More Data
We compare Sundial with other time series foundation models that are pre-trained with smaller datasets: Chronos (ansari2024chronos) is pre-trained on 94 billion time points, and Moirai is pre-trained on 230 billion time points. As their pre-training datasets are part of the subset of TimeBench, we also conduct pre-training on Sundial using these subsets. As shown in Table 8, these results highlight the scaling behavior of Sundial using larger datasets. Additionally, Sundial still achieves better zero-shot forecasting than its counterpart models with the same pre-training dataset.
Table 8: Zero-shot forecasting performance of models trained on different scales of datasets (measured in time points, pts, and 1B means a billion). We report the averaged results from four prediction lengths $\{96,192,336,720\}$ on Time-Series-Library (wu2022timesnet).
@@ C.3 Performance with Varying Lookback Lengths
Figure 10: Zero-shot forecasting performance using different lookback lengths in $\{480,960,1440,1920,2400,2880\}$. We report the averaged results from four prediction lengths $\{96,192,336,720\}$ on Time-Series-Library (wu2022timesnet).
@@ C.4 Zero-Shot Results of Point Forecasting
Table 9 provides full zero-shot results on Time-Series-Library forecasting benchmark (wu2022timesnet), including prediction horizons in $\{96,192,336,720\}$. We build Sundial with different model sizes with configurations in Table 5. The context length is fixed as $2880$. We truncate the model’s predictions for tasks requiring a prediction length less than $F=720$.
We compare the most advanced time series foundation models based on their official checkpoints, including Time-MoE (shi2024time), Timer (liu2024timer; liutimer), Moirai (woo2024unified), TimesFM (das2023decoder), and Chronos (ansari2024chronos). We conduct zero-shot evaluations on datasets that are not included during the pre-training of the corresponding models. For each of the evaluated model, we use their maximum input length during inference. Metrics (MSE/MAE) are calculated from all predicted windows in the test split of each dataset following liu2024timer.
@@ C.5 Zero-Shot Results on GIFT-Eval and FEV Leaderboard
We evaluate our models on GIFT-Eval, a benchmark designed to comprehensively assess forecasting performance across diverse time series. GIFT-Eval includes $23$ datasets covering $144,000$ time series and $177$ million data points, which constitute a total of $97$ forecasting configurations. We use the official evaluation suite established by the research team of Salesforce and report aggregated results in Table 2. We evaluate the performance and inference time on the FEV leaderboard, which was originally proposed by ansari2024chronos and established by AutoGluon, which comprises $27$ datasets for zero-shot evaluation. We report aggregated metrics in Figure 4 and assess the inference time in Figure 5. We released the detailed results by submitting Sundial to their open benchmark [^2] https://huggingface.co/spaces/Salesforce/GIFT-Eval..
@@ D.1 Showcases of Sundial
Figure 11-13 present zero-shot forecasting showcases on all the datasets from FEV (ansari2024chronos) and TSLib (wu2022timesnet). By generating $20$ predictions with different initial noise, we estimate the median and $80\%$ prediction interval.
Table 9: Zero-shot forecasting results of time series foundation models on long-term forecasting datasets (wu2022timesnet). A lower MSE or MAE indicates a better prediction. Averaged results of four prediction lengths are reported here. $1^{\text{st}}$ Count represents the number of wins achieved by a model under all prediction lengths and datasets. Results of baseline models are officially reported by shi2024time. Datasets for pre-training are not evaluated on corresponding models, which are denoted by the dash (-).
@@
$\textbf{Time-MoE}_{\textit{Base}}$
$\textbf{Time-MoE}_{\textit{Large}}$
$\textbf{Time-MoE}_{\textit{Ultra}}$
@@
  • $\ast$ Traffic (trafficdata) is not evaluated because it is included in the pre-training datasets of these time series foundation models.
Figure 11: Showcases of zero-shot predictions from Sundial (Base) on the FEV leaderboard (ansari2024chronos).
Figure 12: Showcases of zero-shot predictions from Sundial (Base) on the FEV leaderboard (ansari2024chronos).
Figure 13: Showcases of zero-shot predictions from Sundial (Base) on long-term forecasting datasets (wu2022timesnet).